How to solve for abnoral return

An index model regression applied to past monthly returns in Ford’s stock price produces the following estimates, which are believed to be stable over time: rF = .10%+ 1.1rM If the market index subsequently rises by 7.3% and Ford’s stock price rises by 7%, what is the abnormal change in Ford’s stock price?
Answer #1
In finance, an abnormal return is the difference between the actual return of a security and the expected return. Abnormal returns are sometimes triggeredby "events." Events can include mergers, dividend announcements, company earning announcements, interest rate increases, lawsuits, etc. all which cancontribute to an abnormal return. Events in finance can typically be classified as occurrences or information that has not already been priced by themarket.

In stock market trading, abnormal returns are the differences between a single stock or portfolio's performance and the expected return over a set periodof time.[1] Usually a broad index, such as the S&P 500 or a national index like the Nikkei 225, is used as a benchmark to determine the expectedreturn. For example if a stock increased by 5% because of some news which affected the stock price, but the average market only increased by 3% and thestock has a beta of 1, then the abnormal return was 2% (5% - 3% = 2%). If the market average performs better (after adjusting for beta) than the individualstock then the abnormal return will be negative.

textrm{Abnormal Return} = textrm{Actual Return} - textrm{Expected Return}
answered by: Herosugar
Answer #2
Abnormal changes are the one that cannot be explained,

now market index rises by 7.3%
according to equation ford's stock should be = .10% + 1.1*7.3 = 8.13%
But it is only 7%

Hence abnormal change = 8.13-7 = 1.13%
answered by: mee
Answer #3
abnormal change in Ford’s stock price= [.10% + 1.1x7.3]-7
= 1.13%
answered by: Bankai